# Deep Smart Contract Intent Detection

### v2

@@ Abstract
In recent years, researchers in the software security field have focused on detecting vulnerabilities in smart contracts to avoid significant losses of crypto assets on the blockchain. Despite early successes in this domain, detecting developers’ intents in smart contracts is a more pressing issue, as malicious intents have resulted in substantial financial losses. Unfortunately, existing research lacks effective methods for detecting development intents in smart contracts.
To address this gap, we propose SmartIntentNN (Smart Contract Intent Neural Network), a deep learning model designed to automatically detect development intent in smart contracts. SmartIntentNN utilizes a pre-trained sentence encoder to generate contextual representations of smart contract code, a K-means clustering model to identify and highlight prominent intent features, and a bidirectional LSTM-based deep neural network for multi-label classification.
We trained and evaluated SmartIntentNN on a dataset comprising over 40,000 real-world smart contracts, employing self-comparison baselines in our experimental setup. The results demonstrate that SmartIntentNN achieves an F1-score of 0.8633 in identifying intents across 10 distinct categories, outperforming all baselines and filling the gap in smart contract detection by incorporating intent analysis.
@@ I Introduction
What are external risks? When considering smart contracts as computer programs, the primary external risks stem from the exploitation of vulnerabilities [11]. A notable example is the infamous DAO attack, triggered by a reentrancy vulnerability [12, 13], in which an attacker repeatedly invoked a function before the previous execution completed, enabling unauthorized withdrawals. Another common issue is unchecked exceptions [14, 15, 16], where improper error handling leads to unexpected contract behavior. Additionally, integer overflow/underflow [17] can cause arithmetic operations to wrap around, leading to incorrect account balances or infinite loops. More vulnerabilities in smart contracts have been comprehensively summarized by Chu et al. [18]. These external risks are attacks from outside the smart contract that exploit its vulnerabilities.
@@
What are internal risks? In legal contracts, malicious terms can lead to significant losses for users. Similarly, in smart contracts, which are transaction protocols designed by developers, harmful terms can be embedded in the form of computer code. Our investigation reveals that, in recent years, an increasing number of risks have been intentionally injected by smart contract developers or DApp creators [19, 20, 21, 22]. Malicious DApps purposely inject tricks and backdoors in their smart contracts to misappropriate users’ funds [23].
@@
Our research focuses on detecting internal risks in smart contracts by identifying developers’ intent embedded within the code. There is already a substantial body of successful research dedicated to analyzing external risks, particularly in smart contract vulnerability detection [18]. However, few studies address internal risks, and no existing methods can detect developers’ intent from the context of smart contract code. Currently, the only way to identify such intents is through expert manual audits, which are time-consuming and costly [24]. This gap in the detection of unsafe developer intents motivates our research.
We present SmartIntentNN, a deep learning model comprising three key components: a Universal Sentence Encoder (USE) [25] for contextual representation of smart contracts, a trained K-means clustering model [26, 27] to highlight distinct intent, and a bidirectional long short-term memory (BiLSTM) network [28, 29] for multi-label classification. SmartIntentNN is implemented using TensorFlow.js [30, 31] and can detect ten categories of intents (see Table I).
We trained and evaluated SmartIntentNN on a dataset of over $40,000$ labeled smart contracts, comparing it against self-comparison baselines, including classic LSTM [28], BiLSTM [29], CNN [32] models, and generative large language models (LLMs) such as GPTs [33]. SmartIntentNN achieved an F1-score of $0.8633$, accuracy of $0.9647$, precision of $0.8873$, and recall of $0.8406$, surpassing all baselines.
Our contributions can be summarized as follows:
  • To the best of our knowledge, this is the first study to propose an approach for detecting developer intents behind smart contracts using deep learning models.
  • We compiled a comprehensive dataset of over $40,000$ labeled smart contracts, covering 10 categories of development intents.
  • We provide a website for SmartIntentNN, with all documentation, code, dataset, and models available at https://github.com/web3se-lab/web3-sekit.
@@ II Motivation
Why is it essential to identify the developers’ intent in smart contract development?For the love of money is the root of all evil. In traditional web software development, the concept of developers’ intent is rarely discussed, as these applications are typically released by a trustworthy and centralized entity. These conventional applications generate revenue by providing high-quality software services to users and do not rely on their cryptocurrency systems. The entities behind these applications are responsible for the code logic and ensuring the software’s overall quality.
In stark contrast, Web3 applications are decentralized and financially driven, allowing any developer to deploy a smart contract and promote it to users for financial purposes. More critically, smart contracts in Web3 applications often include their economic systems, which are tied to real-value cryptocurrencies (e.g., Bitcoin). This is where potential risks emerge: as a smart contract accrues significant value [^2] A way to estimate the value of a smart contract is to calculate its TVL, referring to https://academy.binance.com/en/glossary/total-value-locked-tvl., unscrupulous developers may exploit preset malicious code to steal or misappropriate users’ cryptocurrency.
Given the decentralized nature and the immense financial stakes involved in DApp development, understanding and scrutinizing the intent of developers becomes crucial. Without such scrutiny, users are at a heightened risk of falling victim to malicious activities that can lead to substantial financial losses.
Did developers’ malicious intent cause financial damage? According to the 2022 Crypto Crime Report by Chainalysis [34], cryptocurrency scammers have stolen approximately \$7.8 billion worth of cryptocurrency from victims, with over \$2.8 billion resulting from rug pulls. A rug pull occurs when developers invoke malicious functions embedded in smart contracts to perform unfair transactions, such as illegally withdrawing funds or unpairing liquidity pools to misappropriate crypto assets [35, 36]. Compared to data from 2020, the losses in 2021 surged by \$82%. Further investigation by HoneyBadger revealed that 690 honeypot smart contracts accumulated more than \$90,000 in profit for their creators [19].
Most users are not programming experts and cannot audit the underlying code of these smart contracts, making them easy targets for malicious schemes like rug pulls, honeypots, and phishing scams. In the following sections, we will further detail and categorize the various potential malicious intents we have identified. These malicious actions by developers have caused and will continue to cause irreversible financial damage to users. Therefore, there is an urgent need for effective detection measures to identify and expose these unsafe intents in Web3 development.
@@ III-A Malicious Smart Contract Intent
Smart contracts facilitate online interactions via blockchain networks, and Ethereum, the most widely used smart contract platform, serves as an illustrative example of how developers with malicious intents can exploit these contracts to steal users’ funds. Ethereum and similar blockchain platforms function as transaction-based systems, where both deploying and invoking smart contracts require transactions. Users are represented by externally owned accounts (EOAs), which are necessary for paying gas fees in ether (ETH), Ethereum’s valuable native cryptocurrency.
Figure 2: A description of how developers exploit smart contracts to illegally benefit at the expense of users, detailing the process of creating and deploying contracts with malicious intents.
As illustrated in Fig. 2, both developers and users maintain EOAs. A developer writes a smart contract, compiles it to bytecode executable on the Ethereum Virtual Machine (EVM) [37], and deploys it by signing a transaction with their private key, generating a unique contract address. Malicious developers might incorporate a reward function into the smart contract to lure users, prompting them to stake ETH in exchange for incentives such as NFTs (non-fungible tokens) [38].
A user might send a transaction to invoke a deposit function, paying 10 ETH to the smart contract address. Afterward, the user attempts to invoke the reward function to receive NFTs. However, if the developer has secretly implemented a withdraw function, they can steal the user’s 10 ETH immediately. Once these transactions are confirmed on the blockchain, they cannot be altered or traced back, and the stolen funds are irretrievable.
@@ III-B Sentence Embedding
USE takes a sentence as input and outputs a contextual representation, ensuring that similar sentences are close in the generated vector space [39, 43]. In this paper, we treat every function snippet in a smart contract as a sentence and pass it into USE to obtain the corresponding contextual representation.
@@ IV Dataset
To train and evaluate SmartIntentNN, we compiled an extensive dataset of over 40,000 smart contracts from the Binance Smart Chain (BSC) explorer [^3] https://bscscan.com. BSC is an Ethereum-compatible blockchain, enabling smart contracts on BSC to also operate on Ethereum. Fig. 3 illustrates the data collection and preprocessing workflow.
Initially, we downloaded a substantial number of open-source smart contracts from the blockchain explorer. Then, we merged smart contracts containing multiple files and removed redundant or noisy code snippets. Finally, we employed regular expressions (RegEx) to extract code snippets based on Solidity syntax elements, generating a structured smart contract code tree.
@@ IV-A Intent Labels
The dataset comprises ten categories of unsafe intents, as detailed in Table I. These labeling data were compiled from StaySafu [^4] https://www.staysafu.org, along with the expertise of DApp developers and auditors.
TABLE I: Categories of Risky Intents
@@
Each of these intents can appear multiple times in a smart contract. Our statistical analysis revealed that the most frequent intent is Reflect, appearing in $33,268$ instances out of over $40,000$ smart contracts, accounting for $37.5\%$ of all intents. This is followed by Fee at $26.86\%$ and MaxTX at $13.76\%$. The intent with the least occurrence is MaxSell, appearing in only $68$ instances and accounting for a mere $0.05\%$ of our dataset. These statistics indicate a higher prevalence of certain risks, such as Reflect, Fee, and MaxTX, in smart contracts on the BSC.
@@ V Approach
In this section, we present a DNN-based approach leveraging the SmartIntentNN model for detecting development intents in smart contracts. As illustrated in Fig. 4, SmartIntentNN consists of three core components. First, smart contract embedding using the USE converts the contract context into embedding vectors. Next, the intent highlight model with a K-means model emphasizes features with strong intents. Finally, we feed the highlighted data into a BiLSTM-based DNN layer to learn the representations of smart contracts and perform multi-label classification. A detailed explanation of our approach is provided in the following subsections.
@@ V-A Smart Contract Embedding
To generate contextual representations of a smart contract, we begin by considering each function within it. Each function is individually encoded, and these function embeddings are then combined to create a holistic representation of the smart contract.
@@ V-B Intent Highlight
Feeding $\bm{X}$ directly into a DNN may fail to capture all of the developer’s specific intents in a smart contract. Thus, we introduce an intent highlight model to emphasize intent-related functions. This model employs K-means clustering to predict the intent strength of each vector $\bm{f}$ in $\bm{X}$. By calculating intent strength, we enhance the features of vectors that strongly reflect the developer’s intent. This mechanism is illustrated by the formula $\bm{X^{\prime}}=\mathrm{H}(\bm{X})$, where $\mathrm{H}$ denotes the K-means-based intent highlight model, and $\bm{X^{\prime}}$ is the output of the intent-highlighted data.
To find the appropriate $k$ value for K-means clustering, we compute the occurrence rate of each function $\mathcal{F}$ within a randomly selected subset of the entire smart contract dataset. This subset, $\mathbf{S}$, consists of $m$ smart contracts, denoted as $\mathbf{S}=\left\{\mathbf{T_{i}}\right\}_{i=1}^{m}$. The occurrence rate, $\mathrm{R}\left(\mathcal{F}\right)$, is defined as the frequency with which a function appears in the subset, as detailed in Formula 7. If a function $\mathcal{F}$ exists within the smart contract tree $\mathbf{T_{i}}$, it is counted as $\mathbb{I}\left\{\mathcal{F}\in\mathbf{T_{i}}\right\}$. We specifically tally the number of $\mathcal{F}$ functions whose occurrence rate $\mathrm{R}\left(\mathcal{F}\right)>\rho$, where $\rho$ is an experimentally determined threshold. The total count of these high-frequency functions gives us the value of $k$.
@@
We compute the cosine distance between their embedding vectors when comparing the similarity of two documents [47]. Consider two functions ($\mathcal{F}_{A}$ and $\mathcal{F}_{B}$) with embedding vectors $\bm{f^{A}}=[f_{1}^{A},f_{2}^{A},\dots,f_{n}^{A}]$ and $\bm{f^{B}}=[f_{1}^{B},f_{2}^{B},\dots,f_{n}^{B}]$. Formula 8 calculates the text similarity between $\mathcal{F}_{A}$ and $\mathcal{F}_{B}$, which is the cosine value of $\bm{f^{A}}$ and $\bm{f^{B}}$. We use Formula 9 to transform cosine similarity into cosine distance.
@@
To identify the centroids for K-means clustering, we start with a training dataset $\left\{\bm{f^{\prime}_{i}}\right\}_{i=1}^{n}$, where each $\bm{f^{\prime}_{i}}$ is an embedding vector representing a function $\mathcal{F}$ from the previously described subset $\mathbf{S}$. We initialize $k$ random centroids $\left\{\bm{c_{j}^{t}}\right\}_{j=1}^{k}$ for $t=0,1,\dots,z$, where $z$ is the maximum number of iterations. K-means iteratively updates centroids by minimizing the cosine distance, as opposed to the original Euclidean distance. Each function embedding $\bm{f^{\prime}_{i}}$ is assigned to the closest centroid $\bm{c_{j}^{t}}$, forming a set $\mathbf{M_{j}^{t}}$ for each centroid. The centroid $\bm{c_{j}^{t+1}}$ is updated based on the weighted average of all vectors in $\mathbf{M_{j}^{t}}$, as shown in Formula 10.
@@
The objective is to minimize the total within-cluster variation (TWCV), as depicted in Formula 11. If $\bm{f^{\prime}_{i}}$ belongs to $\mathbf{M_{j}^{t}}$, the cosine distance is calculated. Here, $\mathrm{D}(\bm{f^{\prime}_{i}},\bm{c_{j}^{t}})$ represents the cosine distance between the function vector $\bm{f^{\prime}_{i}}$ and the centroid $\bm{c_{j}^{t}}$ at iteration $t$.
@@
After each iteration, the centroid $\bm{c_{j}^{t}}$ is updated to $\bm{c_{j}^{t+1}}$. The training process continues until $t$ reaches the maximum number of iterations, $z$, or the reduction in TWCV becomes negligible. Ultimately, this process identifies the most appropriate centroids $\left\{\bm{c_{j}}\right\}_{j=1}^{k}$ for K-means.
@@
Once the K-means model has been trained, we can input any function embedding vector $\bm{f_{i}}$ to calculate its within-cluster distance. Around the centroids $\bm{c_{j}}$ in Fig. 5, we observe some frequently occurring and similar function code snippets. Their vectors cluster densely around the centroids. These code snippets typically originate from public libraries, mainstream algorithms, or widely reused functions, indicating weak developer intent. In contrast, centroid-outlying functions are distinctive and express strong intent. Therefore, we consider the within-cluster distance as a measure of intent strength.
@@
We scale $\bm{f_{i}}$ in $\bm{X}$ by the predicted within-cluster distance to generate a new matrix $\bm{X^{\prime}}\in\mathbb{R}^{n\times m}$. In Formula 12, $i\in\left\{1,2,\cdots,n\right\}$ and $j\in\left\{1,2,\cdots,k\right\}$, where $\lambda$ is a distance threshold and $\mu$ is the scaling factor applied to $\bm{f_{i}}$ if its within-cluster distance $\mathrm{D}\left(\bm{f_{i}},\bm{c_{j}}\right)\geq\lambda$. These steps are illustrated in the middle part of Fig. 4 near the intent highlight model.

### v3

@@ Abstract
In recent years, research in software security has concentrated on identifying vulnerabilities in smart contracts to prevent significant losses of crypto assets on blockchains. Despite early successes in this area, detecting developers’ intents in smart contracts has become a more pressing issue, as malicious intents have caused substantial financial losses. Unfortunately, existing research lacks effective methods for detecting development intents in smart contracts.
To address this gap, we propose SmartIntentNN (Smart Contract Intent Neural Network), a deep learning model designed to automatically detect development intents in smart contracts. SmartIntentNN leverages a pre-trained sentence encoder to generate contextual representations of smart contracts, employs a K-means clustering model to identify and highlight prominent intent features, and utilizes a bidirectional LSTM-based deep neural network for multi-label classification.
We trained and evaluated SmartIntentNN on a dataset containing over 40,000 real-world smart contracts, employing self-comparison baselines in our experimental setup. The results show that SmartIntentNN achieves an F1-score of 0.8633 in identifying intents across 10 distinct categories, outperforming all baselines and addressing the gap in smart contract detection by incorporating intent analysis.
@@ I Introduction
What are external risks? When considering smart contracts as computer programs, the primary external risks stem from the exploitation of vulnerabilities [11]. A notable example is the infamous DAO attack, triggered by a reentrancy vulnerability [12, 13], in which an attacker repeatedly invoked a function before the previous execution completed, enabling unauthorized withdrawals. Another common issue is unchecked exceptions [14, 15, 16], wherein improper error handling leads to unexpected contract behavior. Additionally, integer overflow/underflow [17] can cause arithmetic operations to wrap around, leading to incorrect account balances or infinite loops. More vulnerabilities in smart contracts have been comprehensively summarized by Chu et al. [18]. These external risks are attacks from outside the smart contract that exploit its vulnerabilities.
@@
What are internal risks? In legal contracts, malicious terms can lead to significant losses for users. Similarly, in smart contracts, which are transaction protocols designed by developers, harmful terms can be embedded in the form of computer code. Our investigation reveals that, in recent years, an increasing number of risks have been intentionally injected by smart contract developers or DApp creators [19, 20, 21, 22]. Malicious DApps purposely inject tricks and backdoors into their smart contracts to misappropriate users’ funds [23].
@@
Our research focuses on detecting internal risks in smart contracts by identifying negative developer intents embedded within the code. While there is extensive research on external risks, primarily in the area of smart contract vulnerability detection [18], few studies address internal risks. Moreover, no existing methods specifically target the detection of developer intents from the context of smart contract code. Currently, identifying such intents relies on expert manual audits, which are both time-consuming and costly [24].
We present SmartIntentNN, a deep learning model comprising three key components: a Universal Sentence Encoder (USE) [25] for contextual representation of smart contracts, a K-means clustering model [26, 27] to highlight distinctive intent features, and a bidirectional long short-term memory (BiLSTM) network [28, 29] for multi-label classification. Our model is implemented using TensorFlow.js [30, 31] and can detect ten categories of intents (see Table I).
We trained and evaluated SmartIntentNN on a dataset of over 40,000 smart contracts, comparing it against self-comparison baselines, including classic LSTM [28], BiLSTM [29], CNN [32] models, and generative large language models (LLMs) such as GPTs [33]. SmartIntentNN achieved an F1-score of 0.8633, an accuracy of 0.9647, precision of 0.8873, and recall of 0.8406, surpassing all baselines.
Our contributions are summarized as follows:
  • To the best of our knowledge, this study is the first to propose a method for detecting developer intents in smart contracts using deep learning techniques.
  • We have compiled a comprehensive dataset of over 40,000 labeled smart contracts, covering 10 categories of developer intents.
  • We have open-sourced SmartIntentNN, including all documentation, source code, datasets, and models. Please refer to https://github.com/web3se-lab/web3-sekit.
@@ II Motivation
Why is it essential to identify the developers’ intent in smart contract development?For the love of money is the root of all evil. In traditional web software development, the concept of developers’ intent is rarely discussed, as these applications are typically released by a trustworthy and centralized entity. These conventional applications make revenue by providing high-quality software services and do not rely on cryptocurrency systems. The entities behind these applications are responsible for the overall quality of the software.
In stark contrast, Web3 applications are decentralized and financially driven, allowing any developer to deploy a smart contract and promote it to users for financial purposes. More critically, smart contracts in Web3 applications often include their economic systems, which are tied to real-value cryptocurrencies (e.g., Bitcoin). This is where potential risks arise. As the assets held within a smart contract significantly increase in value, unscrupulous developers can exploit embedded malicious code to misappropriate users’ cryptocurrencies.
Given the decentralized and value-driven nature of DApp development, it is crucial to detect and identify developers’ intents. Without such a software process, DApp users are at a heightened risk of exposure to malicious activities, which can result in substantial financial losses.
Did developers’ malicious intent cause financial damage? According to the 2022 Crypto Crime Report by Chainalysis [34], cryptocurrency scammers have stolen approximately \$7.8 billion worth of cryptocurrency from victims, with over \$2.8 billion resulting from rug pulls. A rug pull occurs when developers invoke malicious functions embedded in smart contracts to perform unfair transactions, such as illegally withdrawing funds or removing liquidity pools to misappropriate crypto assets [35, 36]. Compared to data from 2020, the losses in 2021 surged by \$82%. Further investigation by HoneyBadger revealed that 690 honeypot smart contracts accumulated more than \$90,000 in profit for their creators [19].
The malicious intent of developers has indeed caused, and will likely continue to cause, significant financial damage to DApp users. Common schemes such as rug pulls, honeypots, and phishing scams exemplify this risk.
@@ III-A Malicious Smart Contract Intent
Smart contracts facilitate online interactions via blockchain networks, with Ethereum being the most widely used platform. It serves as an illustrative example of how developers with malicious intents can exploit these contracts to steal users’ funds. Ethereum and similar blockchain platforms operate as transaction-based systems, where both deploying and invoking smart contracts require transactions. Users are represented by externally owned accounts (EOAs), necessary for paying gas fees in ether (ETH), Ethereum’s native cryptocurrency.
Figure 2: A depiction of how developers exploit smart contracts for illegal gain, illustrating the process of creating and deploying contracts with malicious intents.
As shown in Fig. 2, both developers and users maintain EOAs. A developer writes a smart contract, compiles it into bytecode executable on the Ethereum Virtual Machine (EVM) [37], and deploys it by signing a transaction with their private key, generating a unique contract address. Malicious developers might incorporate a reward function into the smart contract to lure users, prompting them to stake ETH in exchange for incentives such as NFTs (non-fungible tokens) [38].
A user might send a transaction to invoke a deposit function, paying 10 ETH to the smart contract address. Subsequently, the user attempts to invoke the reward function to receive NFTs. However, if the developer has secretly embedded a withdraw function, they can immediately steal the user’s 10 ETH. Once these transactions are confirmed on the blockchain, they cannot be altered or traced back, rendering the stolen funds irretrievable.
@@ III-B Sentence Embedding
USE takes a sentence as input and outputs a contextual representation, ensuring that similar sentences are close in the generated vector space [39, 43]. In this paper, we treat every function snippet in a smart contract as a sentence and pass it into USE to obtain the contextual representation.
@@ IV Dataset
To train and evaluate SmartIntentNN, we compiled an extensive dataset of over 40,000 smart contracts from the Binance Smart Chain (BSC) explorer [^2] https://bscscan.com. BSC is an Ethereum-like blockchain, enabling smart contracts on BSC to also operate on Ethereum.
We illustrate the data collection and preprocessing workflow in Fig. 3. Initially, we downloaded a number of open-source smart contracts from the blockchain explorer. Then, we merged smart contracts that contained multiple files and cleaned redundant contracts. Finally, we employed regular expressions (RegEx) to extract code snippets based on Solidity syntax keywords, generating a structured smart contract code tree.
@@ IV-A Intent Labels
The dataset includes ten categories of unsafe intents in smart contracts, as detailed in Table I. These ten common intents were collected from the experiences of DApp developers but do not represent all potential categories of intents. The smart contracts in our dataset were labeled by experts, developers, auditors, and through various blockchain security websites, e.g., StaySafu [^3] https://www.staysafu.org.
TABLE I: Categories of Intents
@@
Each of these intents can appear multiple times in one smart contract. Our statistical analysis revealed that the most frequent intent is Reflect, appearing in $33,268$ instances out of our dataset, accounting for $37.5\%$ of all intents. This is followed by Fee at $26.86\%$ and MaxTX at $13.76\%$. The intent with the least occurrence is MaxSell, appearing in only $68$ instances and accounting for a mere $0.05\%$. These statistics indicate a higher prevalence of certain risks, such as Reflect, Fee, and MaxTX, in smart contracts on the BSC.
It is important to note that the unsafe intents in our dataset are not equal to malicious intents. Instead, they represent potential internal risks in smart contracts. Only when developers exploit these intents to initiate unfair transactions do they become true malicious intents.
@@ V Approach
In this section, we introduce the SmartIntentNN model and its approach for detecting intents in smart contracts. As illustrated in Fig. 4, SmartIntentNN comprises three core components. First, the context of smart contracts is encoded into embedding vectors using USE. Next, an intent highlight model with K-means clustering emphasizes features with strong intents. Finally, the highlighted data is fed into a BiLSTM-based DNN layer to capture smart contract representations and perform multi-label classification.
@@ V-A Smart Contract Embedding
To generate contextual representations of a smart contract, we begin by considering each function in it. Each function is individually encoded, and their embeddings are then combined to create a holistic representation of the smart contract.
@@ V-B Intent Highlight
In a smart contract, not all functions express the developer’s intent. To enhance the model’s understanding of these intent features, we introduce an intent highlight model to emphasize intent-related functions. This model utilizes K-means clustering to predict the intent strength of each vector $\bm{f}$ in $\bm{X}$. By quantifying intent strength, we amplify the features of vectors that strongly reflect the developer’s intent. This mechanism is represented by the formula 7, where $\mathrm{H}$ denotes the K-means-based intent highlight model, and $\bm{X^{\prime}}$ is the output of the intent-highlighted data.
$\bm{X^{\prime}}=\mathrm{H}(\bm{X})$ (7)
To find the appropriate $k$ value for K-means clustering, we compute the occurrence rate of each function $\mathcal{F}$ within a randomly selected subset of the entire smart contract dataset. This subset, $\mathbf{S}$, consists of $m$ smart contracts, denoted as $\mathbf{S}=\left\{\mathbf{T_{i}}\right\}_{i=1}^{m}$. The occurrence rate, $\mathrm{R}\left(\mathcal{F}\right)$, is defined as the frequency with which a function appears in the subset, as detailed in Formula 8. If a function $\mathcal{F}$ exists within the smart contract tree $\mathbf{T_{i}}$, it is counted as $\mathbb{I}\left\{\mathcal{F}\in\mathbf{T_{i}}\right\}$. We specifically tally the number of $\mathcal{F}$ functions whose occurrence rate $\mathrm{R}\left(\mathcal{F}\right)>\rho$, where $\rho$ is an experimentally determined threshold. The total count of these high-frequency functions ($\mathrm{R}\left(\mathcal{F}\right)>\rho$) gives us the value of $k$.
@@
We compute the cosine distance between their embedding vectors when comparing the similarity of two documents [47]. Consider two functions ($\mathcal{F}_{A}$ and $\mathcal{F}_{B}$) with embedding vectors $\bm{f^{A}}=[f_{1}^{A},f_{2}^{A},\dots,f_{n}^{A}]$ and $\bm{f^{B}}=[f_{1}^{B},f_{2}^{B},\dots,f_{n}^{B}]$. Formula 9 calculates the text similarity between $\mathcal{F}_{A}$ and $\mathcal{F}_{B}$, which is the cosine value of $\bm{f^{A}}$ and $\bm{f^{B}}$. We use Formula 10 to transform cosine similarity into cosine distance.
@@
To identify the centroids for K-means clustering, we start with a training dataset $\left\{\bm{f^{\prime}_{i}}\right\}_{i=1}^{n}$, where each $\bm{f^{\prime}_{i}}$ is an embedding vector representing a function $\mathcal{F}$ from the previously described subset $\mathbf{S}$. We initialize $k$ random centroids $\left\{\bm{c_{j}^{t}}\right\}_{j=1}^{k}$ for $t=0,1,\dots,z$, where $z$ is the maximum number of iterations. K-means iteratively updates centroids by minimizing the cosine distance, as opposed to the original Euclidean distance. Each function embedding $\bm{f^{\prime}_{i}}$ is assigned to the closest centroid $\bm{c_{j}^{t}}$, forming a set $\mathbf{M_{j}^{t}}$ for each centroid. The centroid $\bm{c_{j}^{t+1}}$ is updated based on the weighted average of all vectors in $\mathbf{M_{j}^{t}}$, as shown in Formula 11.
@@
The objective is to minimize the total within-cluster variation (TWCV), as depicted in Formula 12. If $\bm{f^{\prime}_{i}}$ belongs to $\mathbf{M_{j}^{t}}$, the cosine distance is calculated. Here, $\mathrm{D}(\bm{f^{\prime}_{i}},\bm{c_{j}^{t}})$ represents the cosine distance between the function vector $\bm{f^{\prime}_{i}}$ and the centroid $\bm{c_{j}^{t}}$ at iteration $t$.
@@
After each iteration, the centroid $\bm{c_{j}^{t}}$ is updated to $\bm{c_{j}^{t+1}}$. The training process continues until $t$ reaches the maximum number of iterations, $z$, or the reduction in TWCV becomes negligible. Ultimately, this process identifies the most appropriate centroids $\left\{\bm{c_{j}}\right\}_{j=1}^{k}$ for K-means, representing the spatial cluster centers where frequently occurring contract functions are densely grouped.
@@
Once the K-means model has been trained, we can input any function embedding vector $\bm{f_{i}}$ to calculate its within-cluster distance. As illustrated in Fig. 5, functions that are closer to their centroids typically exhibit weaker developer intent. These are often frequently occurring and similar function code snippets, commonly originating from public libraries, mainstream algorithms, or widely reused functions. In contrast, functions that are further from the centroid, known as centroid-outlying functions, tend to be distinctive and express stronger intent. Here, we utilize the within-cluster distance as a measure to quantify the strength of a function’s intent.
@@
Next, we proceed to effectively highlight the intent features within smart contracts. Using the K-means model, we predict the within-cluster distance for each $\bm{f_{i}}$ in $\bm{X}$ and scale the features of selected functions, resulting in a new matrix $\bm{X^{\prime}}\in\mathbb{R}^{n\times m}$. In Formula 13, $i\in\{1,2,\cdots,n\}$ and $j\in\{1,2,\cdots,k\}$, where $\lambda$ denotes a distance threshold and $\mu$ represents the scaling factor applied to $\bm{f_{i}}$ if its within-cluster distance $\mathrm{D}(\bm{f_{i}},\bm{c_{j}})\geq\lambda$. These transformations are visually represented in the central portion of Fig. 4, adjacent to the intent highlight model.